The solubility of a gas is 4.5 g/l at a pressure of 1.0 atm. At what pressure will there be solubility of 45 g/l


1
Expert's answer
2022-03-24T02:20:02-0400

According to Henry's law:

"\frac{S1}{P1}=\frac{S2}{P2}"


S1-4.5g/l

P1-1.0atm

P2-?

S2-45g/l

"\therefore P2=P1×S2÷S1"

"=\frac{1.0atm\times 45g/l}{4.5g/l}=10atm"